OpenStudy (megannicole51):

The solid whose base is the given region and whose cross-sections perpendicular to the x-axis are squares. I know the answer is V=(e^(2)-1)/(2) but I need to show the work and don't know how...please help:D

OpenStudy (megannicole51):

@phi do u know how to do this one by chance?

OpenStudy (phi):

what is the "given region" ?

OpenStudy (megannicole51):

y=e^x the x-axis, and the lines x=0 and x=1. Find the volume. sorry i forgot to add that part!

OpenStudy (phi):



OpenStudy (phi):

can you finish ?

OpenStudy (phi):

as you move along the x-axis, you get squares that have height h = e^x, and thickness dx. the volume of the this slice is h^2 * dx now integrate these slices from x= 0 to x=1

OpenStudy (megannicole51):